Mathematics 133 – Intensive Calculus for Science, 1 Discussion 2 – Periodicity September 16, 2005

Background

The linear, exponential, and logarithm functions we have studied so far are either always increasing or always decreasing. Not everything in the real world behaves this way! In particular, some processes or phenomena repeat themselves. For instance, some quantity might increase for a while, decrease back to its starting point, then repeat the same pattern over and over again. The length of time before the repetition starts is called the period. Anything that has this property is called periodic.

Discussion Questions

A) What are some examples of periodic things or processes in the real world. Try to think of at least three, and say what the period is in each case.

B) We can also talk about periodic functions in mathematics. What would the graph of a periodic function of x look like. Try to draw one and indicate how the period is showing up in your graph.